A)A block of mass M is pulled over a rough horizontal surface. The block has a coefficient of kinetic friction with the surface

b) using the information from the previous question what is the net work done on the block as it is displaced?
Answer #1

given as shown in fig. Data = Jux=0.12 ↑ Tsino T=420 IM= 8.5kg 33.po fa MAN dasm mg N = mg - Tsino = 8.5*9.81 42 x sin 37.5 =initial (u)= 14 w fiec now by ear of motion as : ² u²t gas 8² 14² + 2x 3.1038x5 To= 5.744 mlece & KE = mo? 2 1 / 3 x 8.5x5.74